# Bridging the performance-gap between target-free and target-based reinforcement learning

### v2

@@ Appendix D List of Hyperparameters
Final learning rate $3\times 10^{-4}$
@@
Number of gradient steps 800 000800\,000
@@
Appendix E Individual Learning Curves
E.1 Deep Q-Network with CNN and LayerNorm
@@
E.2 Deep Q-Network with IMPALA and LayerNorm
@@
E.3 Deep Q-Network with CNN
@@
E.4 Conservative Q-Learning with IMPALA and LayerNorm
@@
E.5 Soft Actor-Critic with SimbaV2 and BatchNorm
Figure 19: Reducing the performance gap in online RL on the 77 DMC Hard tasks with the SimbaV2 architecture and BatchNorm. Left: As opposed to iS-SAC, the target-free approach suffers from a low srank, which indicates a lower representation capability. Middle: The percentage of dormant neurons remains low during training for all methods, not exceeding $7\%$. Right: The target-based approach does not benefit from BatchNorm. This is why it is not included in the experiments with TB-SAC.
Figure 20: Per game training curves of iS-SAC, TF-SAC, and TB-SAC with the SimbaV2 architecture and BatchNorm. iS-SAC consistently performs better than or on par with the target-free approach. Interestingly, iS-SAC even outperforms the target-based approach on the humanoid tasks.

### v3

@@ Appendix E List of Hyperparameters
Final learning rate $1\times 10^{-5}$
@@
Number of gradient steps 600 000600\,000
@@
Appendix F Individual Learning Curves
F.1 Deep $Q$-Network with CNN and LayerNorm
@@
F.2 Deep $Q$-Network with IMPALA and LayerNorm
@@
F.3 Deep $Q$-Network with CNN
@@
F.4 Conservative $Q$-Learning with IMPALA and LayerNorm
@@
F.5 Soft Actor-Critic with SimbaV2 and BatchNorm
Figure 25: Reducing the performance gap in online RL on the 77 DMC Hard tasks with the SimbaV2 architecture and BatchNorm. Left: As opposed to iS-SAC, the target-free approach suffers from a low srank, which indicates a lower representation capability. Middle: The percentage of dormant neurons remains low during training for all methods, not exceeding $7\%$. Right: The target-based approach does not benefit from BatchNorm. This is why it is not included in the experiments with TB-SAC. Importantly, all algorithms use $\ell_{2}$-norm as described in Lee et al. (2025).
Figure 26: Per task training curves of iS-SAC, TF-SAC, and TB-SAC with the SimbaV2 architecture and BatchNorm. iS-SAC consistently performs better than or on par with the target-free approach. Interestingly, iS-SAC even outperforms the target-based approach on the humanoid tasks.